What are the x and y intercepts of the graph 4x - 7y = 14? Can you show me step by step how you solved this?
Found 2 solutions by neatmath, rfer:
Answer by neatmath(302)

It is fairly simple to find the x and y intercepts of a line.

To find the y intercept, just set x equal to zero and solve for y!

4(0)-7y=14

-7y=14

y=-2

Thus the y intercept is -2, or more precisely, the point (0,-2)

You might have guessed by now,

to find the x intercept, just set y equal to zero and solve for x!

4x-7(0)=14

4x=14

x=7/2

Thus the x intercept is 7/2, or more precisely, the point (7/2,0)

The easiest way to find the y intercept is to put the equation into slope intercept form or y=mx+b where m is the slope of the line and b is the y intercept!

Thus,

4x-7y=14

4x-14=7y

7y=4x-14

y=4/7(x)-2

Notice how we get the b to end up equaling -2, which is exactly what we figured out above!

By the same token, we also now know the slope of the line, 4/7.

I hope this helps!

Answer by rfer(16322)

y int when x=0
x int when y=0
y int (0,-2)
x int (3.5,0)